livy:

Jack's pancake recipe calls for 2 teaspoons baking powder and 3 cups of flour. How much baking powder is needed if Jack uses 9 cups of flour? Enter your answer in the box. teaspoons

HaydenwH:

how many more cups of flower did he use?

Convert:

So for this we can create a ratio: 2/3 = x/9 Using this we create an equation and solve. 9*2 = 3(x) 18 = 3x 18/3=3x/3 6 = x She used 6 teaspoons of flour.

Convert:

baking powder* not flour

CripQUEZZ:

You can jus divide to get your answer 9 / 3 = 3 So since you multiply three on one side you jus do the same to the other 2 x 3 = 6 So he will use 6 teaspoons of baking powder